# Towards Understanding Why FixMatch Generalizes Better Than Supervised Learning

### v1

@@ 1 Introduction
Deep learning has made significant strides in various domains, showcasing applications in computer vision and natural language modeling (He et al., 2016; Vaswani et al., 2017; Radford et al., 2018; Dosovitskiy et al., 2020; Ho et al., 2020; Mildenhall et al., 2021; Ouyang et al., 2022; Schick et al., 2023). These advances stem from the scalable supervised learning where simultaneously scaling network size and labeled dataset size often enjoys better performance compared with small-scale network and dataset. Unfortunately, in real-world scenarios, labeled data are often scarce. Accordingly, the performance benefits given by a larger dataset can therefore come at a significant cost, since labeling data often requires human efforts and is very expensive, especially for the scenarios where experts are needed for labeling (Sohn et al., 2020; Ouali et al., 2020; Zhou et al., 2020; Zhang et al., 2021a; Pan et al., 2022).
@@
Despite FixMatch’s practical success, its theoretical foundations have not kept pace with its applications. Specifically, it remains unclear how FixMatch and its SL counterpart perform on deep neural networks, though heavily desired. Moreover, few theoretical studies explore the reasons for the practical superiority in test performance of SSL over SL on networks, let alone FixMatch. Most existing theoretical works (He et al., 2022; Ţifrea et al., 2023) focus on analyzing the over-simplified models, e.g., linear learning models, which, however, differ significantly from the highly nonlinear and non-convex neural networks used in real-world SSL scenarios. Due to this gap, these works cannot well reveal the learning mechanism of SSL like FixMatch on networks. Some of other works (Rigollet, 2007; Van Engelen & Hoos, 2020; Guo et al., 2020) view the model as a black-box function under certain restrictive conditions, and their results do not reveal the dependence on the CNN models which is key for the superiority of FixMatch.
Contributions. To solve these issues, we theoretically justify the superior test performance of FixMatch-like SSL over SL in classification tasks, using FixMatch as a case study. We analyze the semantic learning processes in FixMatch and SL, elucidating their test performance differences and motivating the development of an enhanced FixMatch variant. Key contributions are summarized.
Firstly, we prove that on a three-layered CNN, FixMatch achieves better test accuracy than SL. Specifically, under the widely acknowledged multi-view data assumption (Allen-Zhu & Li, 2023), where multiple/single discriminative features exist in multi/single-view data, FixMatch consistently achieves zero training and test classification errors on both multi-view and single-view data. In contrast, while SL achieves zero test classification error on multi-view data, it suffers up to $50\%$ test error on single-view data, showcasing FixMatch’s superior generalization capacity compared to SL.
@@ 2 Related Works
Pseudo-labeling (Scudder, 1965; McLachlan, 1975) and consistency regularization (Bachman et al., 2014; Sajjadi et al., 2016; Laine & Aila, 2016) are the two important principles responsible for the success of modern deep SSL algorithms (Berthelot et al., 2019b; a; Xie et al., 2020; Zhang et al., 2021a; Xu et al., 2021; Wang et al., 2022b; Chen et al., 2023). FixMatch (Xie et al., 2020), as a remarkable deep SSL algorithm, combines these principles with weak and strong data augmentations, achieving competitive results especially when labeled data is limited. Following FixMatch, several works, e.g., FlexMatch (Zhang et al., 2021a), Dash (Xu et al., 2021), FreeMatch (Wang et al., 2022b), and SoftMatch (Chen et al., 2023), try to improve FixMatch by adopting a flexible confidence threshold rather than the hard and fixed threshold adopted by FixMatch. These modern deep SSL algorithms can achieve remarkable test accuracy even trained with one labeled sample per semantic class (Sohn et al., 2020; Zhang et al., 2021a; Xu et al., 2021; Wang et al., 2022b; Chen et al., 2023).
@@ 3.1 Multi-View Data Distribution
Following Allen-Zhu & Li (2023), we employ the multi-view data assumption that suggests each semantic class possesses multiple distinct features—such as car lights and wheels—that can independently facilitate correct classification. To empirically validate this hypothesis, we utilize Grad-CAM (Selvaraju et al., 2017) to identify class-specific regions within images. As shown in Figure 1, Grad-CAM distinctly highlights separate and non-overlapping regions, such as different parts of a car, that contribute to its recognition. These results corroborate the findings of Allen-Zhu & Li (2023), confirming the presence of multiple independent discriminative features within each semantic class.
@@
Now we introduce the multi-view data assumption in Allen-Zhu & Li (2023), which considers a dataset with $k$ semantic classes. Let each sample pair $(X,y)$ consists of the sample $X$, which is comprised of a set of $P$ patches $\{x_{i}\in\mathbb{R}^{d}\}_{i=1}^{P}$, and $y\in[k]$ as the class label. We assume each class $i$ has two discriminative features, $v_{i,1}$ and $v_{i,2}$ in $\mathbb{R}^{d}$, capable of independently ensuring correct classification. While this analysis focuses on two features per class, the methodology extends to multiple features. Below we define $\mathcal{V}$ as the set of all discriminative features across the $k$ classes:
$\mathcal{V}\!=\!\left\{v_{i,1},v_{i,2}\mid\|v_{i,1}\|_{2}=\ \|v_{i,2}\|_{2}=1,v_{i,l}\perp v_{i^{\prime},l^{\prime}}\text{ if }(i,l)\neq\left(i^{\prime},l^{\prime}\right)\right\}_{i=1}^{k},$ (1)
where the conditions ensure the discrimination of each class and each feature in the data. Accordingly, we define the multi- and singe- view distribution $\mathcal{D}_{m}$ and $\mathcal{D}_{s}$, where data from $\mathcal{D}_{m}$ has two features, and data from $\mathcal{D}_{s}$ have one single feature. Set sparsity parameter $s=\mathrm{polylog}(k)$ and constant $C_{p}$.
@@
For the details of Def. 1, please see Def. 7 in Appendix A. According to the definition, a multi-view sample $(X,y)\in\mathcal{D}_{m}$ has patches with two main features $v_{y,1}$ and $v_{y,2}$ plus some noises, while a single-view sample $(X,y)\in\mathcal{D}_{s}$ has patches with only one primary feature $v_{y,1}$ or $v_{y,2}$ plus noises.
@@ 3.2 FixMatch for Training Neural networks
where $F(X)=(F_{1}(X),\cdots,F_{k}(X))\in\mathbb{R}^{d}$ is defined as
$F_{i}(X)=\sum\nolimits_{r\in[m]}\sum\nolimits_{p\in[P]}\overline{\operatorname{ReLU}}(\langle{w_{i,r}},{x_{p}}\rangle),\ \ (\forall i\in[k]).$ (3)
@@
This three-layer network contains the essential components of neural networks, including linear mapping, activation, and a softmax layer, and thus its analysis provides valuable insights into understanding networks trained using SSL. Notably, many other theoretical works also ultilize shallow networks (e.g., two-layer models) to derive insights into deep networks, as seen in the analyses of Li & Yuan (2017); Arora et al. (2019); Zhang et al. (2021b). Additionally, this setup precisely matches the network architecture in Allen-Zhu & Li (2023), enabling direct comparisons between our FixMatch results and the SL findings in Allen-Zhu & Li (2023) in Sec. 4.2.
@@
For FixMatch-like SSLs, at $t$-th iteration, it has two types of losses: 1) a supervised one $L_{s}^{(t)}$ on labeled data, and 2) an unsupervised one $L_{u}^{(t)}$ on unlabeled data. For $L_{s}^{(t)}$, it is cross-entropy loss on labeled dataset $\mathcal{Z}_{l}$:
@@
where $\alpha(X)$ is a weak augmentation applied to input $X$. In practice, $\alpha(X)$ includes a random horizontal flip and a random crop that retains most region of the image (Sohn et al., 2020; Zhang et al., 2021a) which often do not alter the semantic features. So we treat the weak augmentation as an identity map to simplify our analysis. Additionally, experiments in Appendix K.2 also verify that weak augmentation does not have significant effect on the training of SSL.
For the unsupervised loss $L_{u}^{(t)}(X_{u})$, it feeds a weakly-augmented unlabeled sample $\alpha(X_{u})$ into the network to get the model classification probability $\mathbf{logit}_{i}(F^{(t)},\alpha(X_{u}))\ (i\!\in\![k])$. Next, if the maximal classification probability is highly-confident, i.e., $\max_{i}\mathbf{logit}_{i}(F^{(t)},\alpha(X_{u}))\!\geq\!\mathcal{T}_{t}$ with a confidence threshold $\mathcal{T}_{t}\in(0,1]$, FixMatch-like SSLs use it as the pseudo-label to supervise the corresponding strongly-augmented image $\mathcal{A}(X_{u})$:
@@
For strong augmentation $\mathcal{A}(\cdot)$, it often uses CutOut (DeVries & Taylor, 2017) and RandAugment (Cubuk et al., 2020). CutOut randomly masks a large square region of the input image, potentially removing some semantic features. Experimental results in Appendix K.1 confirm the significant impact of CutOut on image semantics and model performance. RandAugment includes various transformations, e.g., rotation, translation, solarization. Appendix K.1 reveals that those augmentations that may remove data semantics also have a large impact on model performance. Based on these findings, we model the probabilistic feature removal effect of $\mathcal{A}(\cdot)$ for our analysis in Sec. 4.1.
@@ 4.1 Results on Test Performance
As discussed in Sec. 3.1, we assume that the training dataset $\mathcal{Z}$ follows the multi-view distribution $\mathcal{D}$ as defined in Def. 1, with multi-view and single-view sample ratios of $1-\mu$ and $\mu$, respectively. Each class $i\in[k]$ in $\mathcal{Z}$ is associated with two i.i.d. semantic features $v_{i,1}$ and $v_{i,2}$, both of which are capable of predicting label $i$. Multi-view samples contain both semantics, while single-view samples contain only one. For clarity, we denote the labeled multi-view and single-view subsets of $\mathcal{Z}$ as $\mathcal{Z}_{l,m}$ and $\mathcal{Z}_{l,s}$, respectively, and the corresponding unlabeled subsets as $\mathcal{Z}_{u,m}$ and $\mathcal{Z}_{u,s}$. Below, we outline the necessary assumptions on the dataset and model initialization.
@@
(a) The training dataset $\mathcal{Z}$ follows the distribution $\mathcal{D}$, and the size of the unlabeled data satisfies $N_{c}=|\mathcal{Z}_{u,m}\cup\mathcal{Z}_{u,s}|=|\mathcal{Z}_{l,m}\cup\mathcal{Z}_{l,s}|\cdot\mathrm{poly}(k)$.
@@
Assumption 2(a) indicates that number of unlabeled data significantly exceeds that of the labeled data, a common scenario given the lower cost of acquiring unlabeled versus labeled data. The Gaussian initialization in Assumption 2(b) accords with the standard initialization in practice, and is mild. Moreover, we also need assumptions on the strong augmentation $\mathcal{A}(\cdot)$ to formulate the effect of consistency regularization in unsupervised loss (5).
@@
(1) For a single-view image, the sole semantic feature is removed with probability $\pi_{2}$.
(2) For a multi-view image, either of the two features, $v_{i,1}$ or $v_{i,2}$, is removed with probabilities $\pi_{1}\pi_{2}$ and $(1-\pi_{1})\pi_{2}$, respectively. We define strong augmentation $\mathcal{A}(\cdot)$ for multi-view data: for $p\in[P]$,
@@
Suppose Assumptions 2,3 holds. For sufficiently large $k$ and $m\!=\!\mathrm{polylog}(k)$, by setting $\eta\!\leq\!1/\mathrm{poly}(k)$, after running FixMatch for $T\!=\!\mathrm{poly}(k)/\eta$ iterations, we have:
@@
(b) Test performance is good. For any multi-view test sample $(X,y)\sim\mathcal{D}_{m}$ and single-view test sample $(X,y)\sim\mathcal{D}_{s}$, with probability at least $1-e^{-\Omega(\log^{2}k)}$, we have
@@
Theorem 4(a) shows that after $T=\mathrm{poly}(k)/\eta$ training iterations, the network $F^{(T)}$ trained by FixMatch can well fit the training dataset $\mathcal{Z}$, and achieves zero classification error. This is because for any training sample $(X,y)$, the predicted value $F_{y}(X)$ for the true label $y$ consistently exceeds the predictions $F_{j}(X)\ (j\neq y)$ for other class labels, ensuring correct classification. More importantly, Theorem 4(b) establishes that the trained network $F^{(T)}$ can also accurately classify test samples $(X,y)\sim\mathcal{D}_{m}\cup\mathcal{D}_{s}$, validating the generalization performance of FixMatch.

### v2

@@ 1 Introduction
Deep learning has made significant strides in various domains, including computer vision and natural language modeling (He et al., 2016; Vaswani et al., 2017; Radford et al., 2018; Dosovitskiy et al., 2020; Ho et al., 2020; Mildenhall et al., 2021; Ouyang et al., 2022; Schick et al., 2023). These advancements largely stem from scalable supervised learning, where increasing both network size and labeled dataset size typically enhances performance. However, in real-world scenarios, labeled data are often scarce. The benefits of larger datasets come at a high cost, as labeling requires human effort and can be prohibitively expensive, particularly in domains that rely on expert annotation (Sohn et al., 2020; Ouali et al., 2020; Zhou et al., 2020; Zhang et al., 2021a; Pan et al., 2022).
@@
Despite FixMatch’s practical success, its theoretical foundations lag behind its applications. Specifically, it remains unclear how FixMatch and its SL counterpart perform on deep neural networks, despite strong interest. Moreover, few theoretical studies investigate why SSL outperforms SL in test performance on networks, let alone FixMatch. Most existing works (He et al., 2022; Ţifrea et al., 2023) analyze oversimplified models, such as linear learners, which differ significantly from the highly nonlinear and non-convex networks used in real-world SSL. Consequently, these studies fail to capture FixMatch’s learning mechanism. Other works (Rigollet, 2007; Van Engelen & Hoos, 2020; Guo et al., 2020) treat models as black-box functions under restrictive conditions, offering insights that do not account for the CNN dependencies crucial to FixMatch’s superiority.
Contributions. To address these issues, we theoretically justify the superior test performance of FixMatch-like SSL over SL in classification tasks, using FixMatch as a case study. We analyze the semantic feature learning processes in FixMatch and SL, explaining their test performance differences and inspiring an improved FixMatch variant. Our key contributions are summarized as follows:
Firstly, we prove that FixMatch achieves superior test accuracy over SL on a three-layer CNN. Specifically, under the widely acknowledged multi-view data assumption (Allen-Zhu & Li, 2023), where multiple/single semantic features exist in multi/single-view data, FixMatch consistently achieves zero training and test classification errors on both multi-view and single-view data. In contrast, while SL achieves zero test classification error on multi-view data, it suffers up to $50\%$ test error on single-view data, showcasing FixMatch’s superior generalization capacity compared to SL.
@@ 2 Related Works
Pseudo-labeling (Scudder, 1965; McLachlan, 1975) and consistency regularization (Bachman et al., 2014; Sajjadi et al., 2016; Laine & Aila, 2016) are the two important principles responsible for the success of modern deep SSL algorithms (Berthelot et al., 2019b; a; Xie et al., 2020; Zhang et al., 2021a; Xu et al., 2021; Wang et al., 2022b; Chen et al., 2023). FixMatch (Sohn et al., 2020), as a remarkable deep SSL algorithm, combines these principles with weak and strong data augmentations, achieving competitive results especially when labeled data is limited. Following FixMatch, several works, e.g., FlexMatch (Zhang et al., 2021a), Dash (Xu et al., 2021), FreeMatch (Wang et al., 2022b), and SoftMatch (Chen et al., 2023), try to improve FixMatch by adopting a flexible confidence threshold rather than the hard and fixed threshold adopted by FixMatch. These modern deep SSL algorithms can achieve remarkable test accuracy even trained with one labeled sample per semantic class (Sohn et al., 2020; Zhang et al., 2021a; Chen et al., 2023).
@@
Feature Learning.
Prior works on feature learning has shed light on how neural networks learn and represent data (Wen & Li, 2021; 2022; Allen-Zhu & Li, 2022; 2023). For instance, Allen-Zhu & Li (2023) explored how ensemble methods and knowledge distillation enhance generalization, while Wen & Li (2021) explained how contrastive learning captures sparse features and avoids spurious dense ones. Empirically, Park et al. (2018) proposed adversarial dropout to adapt neural networks based on their learning status, improving test performance. Despite these advances, to the best of our knowledge, this work is the first to analyze feature learning of neural networks in semi-supervised settings, with our theory-inspired SA-FixMatch achieving superior generalization performance.
@@ 3.1 Multi-View Data Distribution
Following Allen-Zhu & Li (2023), we adopt the multi-view data assumption, which posits that each semantic class comprises multiple distinct semantic features—such as car lights and wheels—that can independently facilitate correct classification. To empirically validate this, we use Grad-CAM (Selvaraju et al., 2017) to identify class-specific regions in images. As shown in Figure 1, Grad-CAM highlights distinct, non-overlapping regions, such as different parts of a car, that contribute to recognition. These results support the findings of Allen-Zhu & Li (2023), confirming the existence of multiple independent semantic features within each class.
@@
Now we introduce the multi-view data assumption in Allen-Zhu & Li (2023), which considers a dataset with $k$ semantic classes. Let each sample pair $(X,y)$ consists of the sample $X$, which is comprised of a set of $P$ patches $\{x_{p}\in\mathbb{R}^{d}\}_{p=1}^{P}$, and $y\in[k]$ as the class label. We assume each class $i$ has two semantic features, $v_{i,1}$ and $v_{i,2}$ in $\mathbb{R}^{d}$, capable of independently ensuring correct classification. While this analysis focuses on two features per class, the methodology extends to multiple features. Below we define $\mathcal{V}$ as the set of all semantic features across the $k$ classes:
$\mathcal{V}\!=\!\left\{v_{i,1},v_{i,2}\mid\|v_{i,1}\|_{2}=\ \|v_{i,2}\|_{2}=1,v_{i,l}\perp v_{i^{\prime},l^{\prime}}\text{ if }(i,l)\neq\left(i^{\prime},l^{\prime}\right)\right\}_{i=1}^{k}.$ (1)
The conditions ensure the distinction of each class and its semantic features. Accordingly, we define the multi-view and singe-view distributions $\mathcal{D}_{m}$ and $\mathcal{D}_{s}$, where data from $\mathcal{D}_{m}$ has two semantic features, and those from $\mathcal{D}_{s}$ have only one. Set sparsity parameter $s=\mathrm{polylog}(k)$ and constant $C_{p}$.
@@
For the details of Def. 1, please see Def. 7 in Appendix A, and also see more explanations in Appendix L. According to the definition, a multi-view sample $(X,y)\in\mathcal{D}_{m}$ has patches with two semantic features $v_{y,1}$ and $v_{y,2}$ plus some noises, while a single-view sample $(X,y)\in\mathcal{D}_{s}$ has patches with only one semantic feature $v_{y,1}$ or $v_{y,2}$ plus noises.
@@ 3.2 FixMatch for Training Neural networks
where $F(X)=(F_{1}(X),\cdots,F_{k}(X))\in\mathbb{R}^{k}$ is defined as
$F_{i}(X)=\sum\nolimits_{r\in[m]}\sum\nolimits_{p\in[P]}\overline{\operatorname{ReLU}}(\langle{w_{i,r}},{x_{p}}\rangle),\quad\forall i\in[k].$ (3)
@@
This three-layer network, comprising linear mapping, activation, and a softmax layer, captures essential neural network components, offering valuable insights into SSL training. Notably, many theoretical studies also use shallow networks (e.g., two-layer models) to gain insights into deep networks (Li & Yuan, 2017; Arora et al., 2019; Zhang et al., 2021b). Moreover, this setup matches the architecture in Allen-Zhu & Li (2023), enabling direct comparisons between our FixMatch results and their SL findings in Sec. 4.2.
@@
For FixMatch-like SSL, at $t$-th iteration, it has two types of losses: 1) a supervised one $L_{s}^{(t)}$ on labeled data, and 2) an unsupervised one $L_{u}^{(t)}$ on unlabeled data. For $L_{s}^{(t)}$, it is cross-entropy loss on labeled dataset $\mathcal{Z}_{l}$:
@@
where $\alpha(X)$ is a weak augmentation applied to input $X$. In practice, $\alpha(X)$ typically consists of a random horizontal flip and a random crop that retains most region of the image (Sohn et al., 2020; Zhang et al., 2021a), which often do not alter the semantic features. Hence, we treat the weak augmentation as an identity map to simplify our analysis. Additionally, experiments in Appendix K.2 confirm that weak augmentation has minimal impact on SSL training.
For the unsupervised loss $L_{u}^{(t)}(X_{u})$, a weakly-augmented unlabeled sample $\alpha(X_{u})$ is fed into the network to obtain classification probabilities $\mathbf{logit}_{i}(F^{(t)},\alpha(X_{u}))$ for $i\in[k]$. If the maximal probability exceeds a confidence threshold $\mathcal{T}_{t}\in(0,1]$, i.e., $\max_{i}\mathbf{logit}_{i}(F^{(t)},\alpha(X_{u}))\geq\mathcal{T}_{t}$, FixMatch-like SSLs use it as the pseudo-label to supervise the corresponding strongly augmented sample $\mathcal{A}(X_{u})$:
@@
For strong augmentation $\mathcal{A}(\cdot)$, it often uses CutOut (DeVries & Taylor, 2017) and RandAugment (Cubuk et al., 2020). CutOut randomly masks a large square region of the input image, potentially removing partial semantic features. Experimental results in Appendix K.1 confirm the significant impact of CutOut on image semantics and model performance. RandAugment includes various transformations, e.g., rotation, translation, solarization. Appendix K.1 reveals that those augmentations that may remove data semantics also have a large impact on model performance. Based on these findings, we model the probabilistic feature removal effect of $\mathcal{A}(\cdot)$ for our analysis in Sec. 4.1.
@@ 4.1 Results on Test Performance
As discussed in Sec. 3.1, we assume the training dataset $\mathcal{Z}$ follows the multi-view distribution $\mathcal{D}$ (Def. 1), with multi-view and single-view sample ratios of $1-\mu$ and $\mu$, respectively. Each class $i\in[k]$ has two i.i.d. semantic features $v_{i,1}$ and $v_{i,2}$, both of which are capable of predicting label $i$. Multi-view samples contain both features, while single-view samples have only one. For clarity, we denote the labeled multi-view and single-view subsets as $\mathcal{Z}_{l,m}$ and $\mathcal{Z}_{l,s}$, and the unlabeled subsets as $\mathcal{Z}_{u,m}$ and $\mathcal{Z}_{u,s}$. Below, we outline the necessary assumptions on the dataset and model initialization.
@@
(a) The training dataset $\mathcal{Z}$ follows the distribution $\mathcal{D}$, and the size of the unlabeled data satisfies $N_{u}=|\mathcal{Z}_{u,m}\cup\mathcal{Z}_{u,s}|=|\mathcal{Z}_{l,m}\cup\mathcal{Z}_{l,s}|\cdot\mathrm{poly}(k)$.
@@
Assumption 2(a) indicates that number of unlabeled data significantly exceeds that of the labeled data, a common scenario given the lower cost of acquiring unlabeled versus labeled data. The Gaussian initialization in Assumption 2(b) accords with the standard initialization in practice, and is mild. Moreover, we also need assumptions on the strong augmentation $\mathcal{A}(\cdot)$ to formulate the effect of consistency regularization in unsupervised loss Eq. (5).
@@
1) For a single-view image, the sole semantic feature is removed with probability $\pi_{2}$.
2) For a multi-view image, either of the two features, $v_{i,1}$ or $v_{i,2}$, is removed with probabilities $\pi_{1}\pi_{2}$ and $(1-\pi_{1})\pi_{2}$, respectively. We define strong augmentation $\mathcal{A}(\cdot)$ for multi-view data: for $p\in[P]$,
@@
Suppose Assumptions 2, 3 hold. For sufficiently large $k$ and $m=\mathrm{polylog}(k)$, setting $\eta\leq 1/\mathrm{poly}(k)$ and running FixMatch for $T=\mathrm{poly}(k)/\eta$ iterations ensures:
@@
(b) Test performance is good. With probability at least $1-e^{-\Omega(\log^{2}k)}$ over the selection of any multi-view test sample $(X,y)\sim\mathcal{D}_{m}$ and single-view test sample $(X,y)\sim\mathcal{D}_{s}$, we have
@@
Theorem 4(a) shows that after $T=\mathrm{poly}(k)/\eta$ training iterations, the network $F^{(T)}$ trained by FixMatch can well fit the training dataset $\mathcal{Z}$, achieving zero classification error. Specifically, for any training sample $(X,y)\in\mathcal{Z}$, the predicted value $F_{y}(X)$ for the true label $y$ consistently exceeds the predictions $F_{j}(X)$ for all $j\neq y$, ensuring correct classification. More importantly, Theorem 4(b) establishes that the trained network $F^{(T)}$ can also accurately classify test samples $(X,y)\sim\mathcal{D}_{m}\cup\mathcal{D}_{s}$, validating FixMatch’s strong generalization performance.